OpenStudy (anonymous):

What is the y-intercept of the equation above? y - 4x = 5

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

slope is 4

OpenStudy (anonymous):

how did you get -5

OpenStudy (johnweldon1993):

y = mx+b...this is called slope intercept form...why? because it shows the slope and y intercept...... m is the slope......b is the y-intercept turn yours into this form y - 4x = 5 y = 4x + 5 now y=mx + b y=4x + 5 m is the slope......b is the y-intercept...so what is your slope...and what is your y-intercept?

OpenStudy (anonymous):

y-4x= 5 -1y -1y -4x=-1y +5 -4x= -1y+5 -5 -5 -5-4x= -1y -5-4x=y -5/-1 -4x/-1 = y 4x-5=y

OpenStudy (anonymous):

So if u know the whole y=mx+b thing this is the way you'd go:) Do u get it? @wild

OpenStudy (anonymous):

yeah but what is the y-intercept

OpenStudy (anonymous):

the y-intercept is -5

OpenStudy (tkhunny):

Whether you know some form or not, the y-intercept corresponds with x = 0. Substitute x = 0 and discover that the y-intercept is (0,5).

OpenStudy (johnweldon1993):

@wild look at my example and follow it...you'll see it

OpenStudy (johnweldon1993):

@LizzyLove<3 "-5-4x= -1y becomes -5-4x=y" when you divide by a -1...you change the signs of both -5 and -4 -5/-1? is +5 -4/-1? is +4

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks i got the answer